In a certain region an electric potential V is present. An unknown charge Q is moved around this region between points at potential difference ΔV. A physicist measures the change in the potential energy ΔU as the charge is moved. The table below shows the data of the measurements. The unit for ΔV are MV (mega-Volts), the units for ΔU are kJ (kilo-joules).

ΔV

57.1

29.9

16.7

9.97

8.55

8.32

6.52

6.87

ΔU

362.8

157.6

124.3

55.3

48.8

53.9

37.9

47.3

You want to find the value of the unknown charge Q in mC. Since ΔU = QΔV make the plot 𝑦=𝑚𝑥 where 𝑦=𝛥𝑈 and 𝑥=𝛥𝑉. Use Excel, calculator or whatever you prefer. From the slope of the plot, calculate Q.


1
Expert's answer
2022-03-11T08:29:30-0500

The equation is the following:


"\Delta U = Q\Delta V"

where "Q" is the charge of the particle. Thus, the graph will be the straight line with slope "Q". Using Excel and the linear approximation, obtain the slope.



Let's account for units:


"[Q] = \dfrac{[\Delta U]}{[\Delta V]} = \dfrac{10^3J}{10^6V} = 10\times 10^{-3}C"

Thus, obtain:


"Q = 6.19\times 10^{-3}C"